OpenStudy (anonymous):

7x+5y=-3 -9x+y=-11 solve linear system.

OpenStudy (help!!!!):

x or y?

OpenStudy (anonymous):

its just like you have to find x and y and it works on both equation using the same x and y

OpenStudy (anonymous):

both x and y

OpenStudy (anonymous):

7x + 5y = -3 -9x + y = -11 (multiply this equation by 5 so the y's are the same) so it 7x + 5y = -3 -45x + 5y = -55 then subtract the two equations to get rid of the y's so it becomes, 52x = -3 -(-55) 52x = 52 x = 1 sub x in any of the two equations to find y, |so 7(1) + 5y = -3 5y= -10 y = -2

OpenStudy (anonymous):

o_o its confusing~~what if its a fraction 3/4x+1/4y=13/2 x-3/4y=13/2

OpenStudy (anonymous):

multiply to get rid of the fraction

OpenStudy (anonymous):

lowest common denominator, so in this case, multiply by 4 to get rid of the denominator

OpenStudy (anonymous):

SO YOU TIMES 4 BY ALL THE #

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so the first equation becomes,3x + y = 26

OpenStudy (anonymous):

write out my answer for the first one, it should help

OpenStudy (anonymous):

3x+1=26 4x-3y=26

OpenStudy (anonymous):

then~~what do i do?

OpenStudy (anonymous):

do the same thing i did above, multiply the first equation by 3 and then add them to get rid of the y's. note it shud be 3x + y = 26

OpenStudy (anonymous):

i got 3x + y = 26 and 4x-3y=26 by multiplying by 4..then i multiply what?

OpenStudy (anonymous):

do the same thing i did above, multiply the first equation by 3 and then add them to get rid of the y's. note it shud be 3x + y = 26

OpenStudy (anonymous):

you want to have the same number of x's or y's so you can eliminate one of them by adding or subtracting.

OpenStudy (anonymous):

in order to solve for the other variable.

OpenStudy (anonymous):

you mulitiply 3x+y=26 or the original equation by 3?